Find the Tangent line of the curve.
xy+4x-3y=0 at the point p(1,2).

Answer by Alan3354(69443)

Find the Tangent line of the curve.
xy+4x-3y=0 at the point p(1,2).
----------
xy+4x-3y=0
Differentiate implicitly
xdy + ydx + 4dx - 3dy = 0
dy*(x-3) + dx*(y+4) = 0
dy/dx = -(y+4)/(x-3)
-----
@ (1,2) y' = -6/(-2)
= 3
-------------
You could solve for y and get the same result.
y = 4x/(3-x)